Hercules Films is deciding on the price of the video release of its film Bride of the Son of Frankenstein. Marketing estimates that at a price of p dollars, it can sell
q = 240,000 − 15,000p
copies, but each copy costs $4 to make. What price will give the greatest profit?
Answer #1

Profit =S(p)= q*p - 4q = 240000p - 15000p2 + 60000p - 960000 = 300000p - 15000p2 - 960000

S'(p) = 300000 - 30000p = 0 -> p = 10

So the price that gives the greatest profit is $10
Answer #2

Given the information:

q= 240000 - 15000p, where C = 4 each copy

So:

Profit =S(p)= q*p - 4q

Profit = 240000p - 15000p^{2} + 60000p - 960000

Profit = = 300000p - 15000p^{2} - 960000

S'(p) = 300000 - 30000p = 0 \rightharpoonup p = 10

Finally, the price that gives the greatest profit is $10.

Let me know if you have any question, and please don't forget rate my answer.